The question is:
Find the original price of a pair of boots if the sales price is $120 after a 20% discount.
I know the original price is $150 because I took 20% and subtracted it from 100% to get 80% and then turned 80 into .80 and divided it into 120 to get the original price but what is the correct actually formula I should have used in order to get the price?

Answer by Maths68(1474)

Original Price = o = ?
Discount = d = 20% of o = (o)*20/100=(o)*1/5=o/5
Sale-price = s = $120
Original Price = Sale-price + discount
o=s+d
o=120+(o/5)
Multiply by 5 both sides of above equation
5*o=5*(120)+0
5o=600+0
5o-o=600
4o=600
divide by 4 both sides of the equation
4o/=600/4
o=150
Original Price = o = $150